If P+1 = N^2 , then find out the value of P !
Here , P= prime number & N = Normal number

Answer by oscargut(2103)

Solution:
P+1 = N^2
P = N^2-1
P = (N-1)(N+1)
But P is prime then N-1=1 or N+1 = 1
N-1 = 1 then N=2 and P = 3
N+1 = 1 then N =0 and P = -1 (not prime)
Answer: P = 3
You can send me more problems to my e-mail: mthman@gmail.com